### v2

# Generalized Sato-Tate and quadratic residues

Generalized Sato-Tate and quadratic residues
Sergey Vlăduţ Aix–Marseille Université, CNRS, I2M UMR 7373, Marseille, France, sergevladuts@ya.ru, vladut@iml.univ-mrs.fr
Aix–Marseille Université, CNRS, I2M UMR 7373, Marseille, France, sergevladuts@ya.ru, vladut@iml.univ-mrs.fr
1. Introduction
The present note is a continuation of [KTVZ] and deals with several questions concerning the distribution of consecutive quadratic residues that were left open there.
One of the main results of [KTVZ] provides rather precise information on the distribution of consecutive quadratic residues modulo very large primes. These results are based on strong statistical independence assumptions that appear natural but are, at present, completely out of reach.
The main observation in this note is that these assumptions follow from certain particular cases of the Generalized Sato–Tate conjecture (GST). Moreover, some already established cases of the GST yield interesting upper and lower bounds for the corresponding arithmetic functions.
Section 2 recalls the formulation of the GST. Section 3 summarizes several results from [KTVZ] concerning consecutive quadratic residues and their connections with algebraic curves and abelian varieties over finite fields. Section 4 presents asymptotic results on the distribution of four and five consecutive quadratic residues, under the assumption of the GST for certain products of elliptic curves of dimension at most five. Finally, Section 5 establishes unconditional bounds, similar in form but weaker than those of Section 4, by using known cases of the GST for products of two elliptic curves.
Acknowledgements. I would like to thank Andrew Sutherland for his very useful remark [Su1]. Many thanks to David Kohel for his help with numerous calculations, especially concerning the curve $C$ on page 8.
2. The Generalized Sato-Tate conjecture
Let us recall the setting for the (generalized) Sato-Tate conjecture and some known elementary cases. Recall that we consider only the simplest case of the base field $\mathbb{Q}.$
Thus, let $A$ be an abelian variety over $\mathbb{Q}$ of dimension $g\geq 1$. The Generalized Sato-Tate conjecture (GST, or GST(A)):
[…]
II. states that the family $\{\mathrm{F}_{p},p\notin S_{A}\}\subset\mathrm{ST}_{A},$ where $p$ is a (rational) prime, formed by natural images of $p$-Frobenius elements $\mathrm{Fr}_{p}\in\mathrm{G}_{\mathbb{Q}}:=\mathrm{Gal}({\bar{\mathbb{Q}}}/\mathbb{Q})$ in $\mathrm{ST}_{A}$ is equidistributed with respect to the probabilistic Haar measure $\mu_{A}$ on $\mathrm{ST}_{A},\;S_{A}$ being the finite set of bad reduction primes for $A.$
We recall then the construction of $\mathrm{F}_{p}$ and the notion of being equidistributed.
Fix a prime $\ell\neq p$ and let
$\phi_{\ell}:\mathrm{G}_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2g}(\mathbb{Q}_{\ell})\subset\mathrm{GL}_{2g}(\bar{\mathbb{Q}}_{\ell})$
be the standard $\ell$-adic representation, given by the Galois action on torsion points of $A.$ Choose and fix an isomorphism $I:\bar{\mathbb{Q}}_{\ell}\longrightarrow\mathbb{C},$ and hence an isomorphism
$I_{2g}:\mathrm{GL}_{2g}(\bar{\mathbb{Q}}_{\ell})\longrightarrow\mathrm{GL}_{2g}(\mathbb{C}),$
then by Weil-Deligne we have
$\mathrm{F}_{p}:=I_{2g}(\phi_{\ell}(\mathrm{Fr}_{p})/\sqrt{p})\in\mathrm{Sp}_{2g}(\mathbb{C})$
and, in fact, $\{\mathrm{F}_{p},p\notin S_{A}\}\subset\mathrm{ST}_{A}$ (and Zarisky-generate $\mathrm{ST}_{A}$).
Let then $T$ be a compact and $\mu:=\mu_{T}$ be a probabilistic Borel measure on $T.$ If $S=\{t_{1},t_{2},\ldots t_{n},\ldots\}\subset T$ then $S$ is equidistributed with respect to $\mu$, or $\mu$-equidistributed if the sequence $\{\lambda_{n}\}$ of discrete measures
$\lambda_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{i}$
*-weakly converges to $\mu$ for $n$ going to infinity, $\delta_{i}$ being the Dirac measure supported on $t_{i}.$ Recall that it means that for any continuous function $f:T\longrightarrow\mathbb{C}$ one has
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f(t_{i})=\int_{T}fd\mu\,.$
Since one has the canonical surjection
$\Phi:\mathrm{Sp}_{2g}(\mathbb{C})\longrightarrow[-2g,2g],M\mapsto\mathrm{Tr}(M),$
we can consider the projected measure $\nu_{A}:=\Phi_{*}(\mu_{A})$ on [2g,2g],[-2g,2g], and the sequence or traces $\{\Phi(\mathrm{F}_{p}),p\notin S_{A}\}$ is $\nu_{A}$-equidistributed.
Example 1. Let $g=1$, i.e. $A=E,$ an elliptic curve over $\mathbb{Q}$ and $\mathrm{ST}_{E}\subset\mathrm{Sp}_{2}(\mathbb{C})=\mathrm{SU}_{2}(\mathbb{C}).$ Then there are two possibilities:
a). $E$ is a CM-curve, that is $\mathrm{End}_{\mathbb{C}}(E)\neq\mathbb{Z};$ then $\mathrm{ST}_{E}=N(T),$ the normalizer of the diagonal subgroup $T$ which has index 2 in $N(T).$
b). $\mathrm{End}_{\mathbb{C}}(E)=\mathbb{Z}$, that is, $E$ has no complex multiplication, and $\mathrm{ST}_{E}=\mathrm{SU}_{2}(\mathbb{C}).$
Note that in any case $\mathrm{End}_{\mathbb{Q}}(E)=\mathbb{Z}.$
The case a) is well-known (Hecke,Weil,…) since sixties, while the equidistribution in the case b) is just the initial Sato-Tate conjecture for elliptic curves, proved in [HSBT].
A simple and well-known calculation (see, e.g., [Se], p.108) shows that
(1) $\nu_{1}=\nu_{CM}=\frac{1}{2}\delta_{0}+\frac{dt}{2\pi\sqrt{4-t^{2}}},\quad\nu_{2}=\nu_{nCM}=\frac{\sqrt{4-t^{2}}dt}{2\pi}.$
In the present note we need only very special cases of GST when $A$ is a product of elliptic curves.
In particular, for $g=\dim A=2$, we use $A=E_{1}\times E_{2}$ for 2 non-isogenous EC $E_{1},E_{2};$ note that in this case GST(A) is known thanks to [Jo].
For $g=3$ we need $A=E_{0}\times E_{1}\times E_{2},$ where $E_{0}$ has CM, and $E_{1},E_{2}$, being non-isogenous (over $\mathbb{C}$) are without CM. This case is not described in the literature, but it could be accessible [Su1] using the methods of [Jo].
However, the next case of interest for us has $g=17$ and thus is far from the current proof methods.
Nevertheless, in the rest of our note we suppose, except in Section 5 that GST(A) is true for all abelian varieties $A$ used below.
3. Consecutive quadratic residues and algebraic curves
Let $p$ be an odd prime. Consider the sequence $1$, 22, …, p1p-1. Replace every number $i$ in the sequence by the letter $R$ if $i$ is a quadratic residue modulo $p$, and by the letter $N$ otherwise. Denote by $W_{p}$ the resulting word.
Definition 1.
Let $S$ be a word of length $t\leq{p-1}$ that contains only $R$ and $N$. Define $n_{p}(S)$ as the number of sub-words of $W_{p}$ that coincide with $S$, and are formed by $t$ consecutive elements of $W_{p}$. The word
$\underbrace{R\ldots R}_{t}$
will be also denoted by $R^{t}$.
Below we will concentrate only on the behaviour of $n_{p}(R^{t})$ and denote it by $n_{p}(t)$. For other cases and some detail see our paper [KTVZ].
Note that if $t=1$, then $n_{p}(1)=\frac{p-1}{2}$, i.e., the number of quadratic residues is equal to the number of non-residues.
Aldanov’s paper [A] gives the answer for $t=2:$
$n_{p}(2)=\frac{p-5}{4},\hbox{ if }p=4k+1,\;n_{p}(2)=\frac{p-3}{4},\hbox{ if }p=4k+3.$
for all (odd) pp.
Jacobstahl’s 1906 thesis [J] explores the case $t=3$, which is trickier. However, for $p=4k+3$, the answer is simple:
[…]
On the contrary, for $p=4k+1$ the formula is not so simple. Namely, define $J(k)\in 2\mathbb{Z}$ by
[…]
Then
[…]
For an arbitrary $t\geq 1,$ and an odd $p$ one can prove [C, p.3], see also [KTVZ, section 2.3]:
[…]
which generalizes Jacobsthal’s formula.
Then a simple calculation, see [KTVZ, section 2.3] gives:
$n_{p}(t)=2^{-t}p+2^{-t}\sum_{s=1}^{t}\sum_{\begin{subarray}{c}T\subseteq[1,t]\\ |T|=s\end{subarray}}N_{T}+c_{p}$
for a certain constant $c_{p}=c_{p}(t)\in 2^{-t}\mathbb{Z}$ which we will mostly ignore. Here
[…]
for
$f_{T}(X):=\prod_{(i+1)\in T}(X+i)\in\mathbb{Z}[X].$
If $C_{T}$ is a hyper-elliptic curve
$C_{T}:y^{2}=f_{T}(x),\quad,$
of genus $g_{T}=g(C_{T})=\lfloor(s-1)/2\rfloor$ for $s=|T|,$ then $p+1-N_{T}$ is just the trace of Frobenius for $C_{T}$.
Note also that for $t\geq 2$ one has
$g_{t}:=\sum_{\begin{subarray}{c}T\subseteq[1,t]\\ \end{subarray}}g_{T}=2^{t-2}(t-3)+1,$
see [KTVZ, section 2.3]
Let us look now at small values of $t$.
$\bf{t=1}.$ Here there is only one $T$ with $f_{T}=X,$ $C_{T}$ is rational, $N_{T}=0$ and $c_{p}=-\frac{1}{2}.$
$\bf{t=2}.$ Here $g_{2}=0,$ there are 3 possible $T,$ all with rational $C_{T},\,N_{T}=0$ and $c_{p}=-\frac{1}{4}$ or $c_{p}=-\frac{3}{4}$ depending on $p\mod 4.$
$\bf{t=3}.$ Here $g_{3}=1;$ there are 7 possible $T,$ all of them, except for $\tilde{T}=\{1,2,3\}$ having rational $C_{T},\,N_{T}=0,T\neq\tilde{T}$; while $C_{\tilde{T}}$ is the ”Gauss” CM elliptic curve $E_{0}$
$y^{2}=x(x+1)(x+2),\;y^{2}=z^{3}-z,\;z:=x+1$
which leads to Jacobsthal’s formula with $c_{p},|c_{p}|\leq\frac{3}{2}$ depending on $p\mod 8.$
tt=4. Here $g_{4}=5$ and we get 5 non-rational curves, all elliptic (4 for $|T|=3$ and 1 for $|T|=4$):
(3) $E_{0}:y^{2}=x(x+1)(x+2);\quad E_{1}:\;y^{2}=x(x+1)(x+3);\quad E_{2}:\;y^{2}=x(x+2)(x+3);$
(4) $E_{3}:\;y^{2}=(x+1)(x+2)(x+3);\quad E_{4}:\;y^{2}=x(x+1)(x+2)(x+3);$
$E_{0}$ and $E_{3}$ are isomorphic over $\mathbb{Q}$, $E_{1},E_{2}$ are isomorphic over $\mathbb{Q}(i)$, but not over $\mathbb{Q}$, whereas $E_{0},E_{3}$ are CM curves, and $E_{1},E_{2},E_{4}$ are not. Therefore,
(5) $n_{p}(4)=\frac{p}{16}+\frac{a_{4}(p)}{16}+c_{p}(4)\;\mbox{for}\;p=4k+3,$
(6) $n_{p}(4)=\frac{p}{16}+\frac{a_{0}(p)}{8}+\frac{a_{1}(p)}{8}+\frac{a_{4}(p)}{16}+c_{p}(4)\;\mbox{for}\;p=4k+1,$
[…]
$\bf{t=5}.$ For $t=5,\;g_{5}=17$ the same approach yields 15 elliptic curves and a curve of genus $2$, namely:
$C_{\{1,2,3,4,5\}}:\;y^{2}=x(x+1)(x+2)(x+3)(x+4).$
However, counting points on this curve can be reduced to counting points on elliptic curves, since this curve has a non-hyperelliptic involution (see below) and its Jacobian splits (over the base field for $p\equiv 1\mod 4$ and over its quadratic extension for $p\equiv 3\mod 4$). Hence, $n_{p}(5)$ can still be expressed in terms of Frobenius traces for elliptic curves.
$\bf{t=6}.$ If $t=6,\;g_{6}=49$, then some genus 2 curves $C_{T}$ for $|T|=5$ have simple Jacobians, which leads to an expression of $n_{p}(6)$ in terms of Frobenius traces for numerous ($\geq 35$) elliptic curves and few ($\leq 7$) curves of genus 2.
$\bf{t=7}$. For $t=7,\,g_{7}=129$, the formulas are of the same type as for $t=6,$ since the only genus 3 curve
$C_{T},|T|=7,y^{2}=x(x+1)(x+2)(x+3)(x+4)(x+5)(x+6)$
has a non-hyperelliptic involution and its Jacobian splits into a product of an elliptic curve and the Jacobian of a curve of genus 2.
Note that for $t\leq 3$ we have ”explicit” formulas for $n_{p}(t)$, but, as explained in [KTVZ, section 2.3], there is no hope to have similar formulas for $t\geq 4$.
4. Asymptotic behaviour
We are now going to show that one can obtain some information on the asymptotic distribution of $n_{p}(t),p\longrightarrow\infty$ from GST, even if no explicit formula for $n_{p}(t)$ is expected for $t\geq 4$. Let us fix $t\geq 1.$
Therefore, the behaviour of $n_{p}(t)$ is governed by the following abelian variety AA
(7) $A=A(t)=\prod_{T\subset[1,...,t]}J_{T},\;J_{T}=\mathrm{Jac}(C_{T}),$
which is the product of all the Jacobians of the curves $C_{T},$ and thus
$\dim(A(t))=g_{t}=2^{t-2}(t-3)+1.$
We are interested in the asymptotic distribution (if it exists) of the quantity
$\delta^{\prime}_{p}(t):=n_{p}(t)-2^{-t}p=\mathrm{trace_{A}(Fr_{p})}+o_{p}(1)=\sum_{T\subset[1,...,t]}\mathrm{trace_{T}(Fr_{p})}+o_{p}(1)$
$\mathrm{trace_{T}(Fr_{p})}$ being the Frobenius trace on $C_{T},$ or rather its normalized version
$\delta_{p}(t)=\delta^{\prime}_{p}(t)/\sqrt{p}\in[-2g_{t},2g_{t}].$
We begin with
4.1. Asymptotics under GS
As mentioned in section 2, we suppose now that GST holds for abelian varieties of the form (7). Therefore, we get that
$\mathrm{ST(A)}\subset\prod_{T\subset[1,...,t]}\mathrm{ST(J_{T})}$
Note, however, that this inclusion is always proper for any $t\geq 3,$ and thus it is not possible to deduce directly from GST the asymptotics of the quantity $\delta_{p}(t).$ Nevertheless, for some small $t$ it is possible to refine the analysis and get some equidistribution results for $\delta_{p}(t);\;p\longrightarrow\infty.$
The first non-trivial case is $t=3,\;g_{t}=1$ which leads to the equidistribution of $\delta_{p}(3)$ under $\nu_{CM},$ and this is unconditional.
4.2. Case $t=4$
For $t=4,\;g_{t}=5$ and $A(t)$ takes the form (see (3),(4))
$A(4)=E_{0}\times E_{1}\times E_{2}\times E_{3}\times E_{4}\equiv E_{0}^{2}\times E_{1}\times\bar{E}_{1}\times E_{4}$
$\bar{E}_{1}$ being the non-trivial twist of $E_{1}$ over $\mathbb{Q}(i)$.
Assuming GST for $B(4):=E_{0}\times E_{1}\times E_{4}$, we get
$\mathrm{ST(B(4))}\subseteq\Delta_{1}\times N\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{\bf{T}}\times\mathrm{SU_{2}(\mathbb{C})}\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{SU_{6}(\mathbb{C})},$
where $\bf{T}=\mathrm{U_{1}(\mathbb{C})}\subset\mathrm{SU_{2}(\mathbb{C})},$ $\Delta_{1}$ being the image of the diagonal embedding $\bf{T}\hookrightarrow\mathrm{SU_{2}(\mathbb{C})}$, whereas $N$ is the normalizer of $\bf{T}\hookrightarrow\mathrm{SU_{2}(\mathbb{C})}$. In (5) and (6) the summands $0,0,a_{4}(p)/16$ and $a_{0}(p)/8,a_{1}(p)/8,a_{4}(p)/16,$ respectively, correspond to the measures
(8) $\mu_{\Delta_{1}}=\nu_{1}(\cdot/2)/2,\;\mu_{N}=\delta_{0}+\nu_{2}/2,\;\nu_{2}.$
Theorem 4.1.
$(i)$ For p running the set $\mathcal{P}_{3}:=\{p=4m+3\;\;is\;\ prime\}$ the quantity $\delta_{p}(4)$ is equidistributed on [2,2][-2,2] with respect to the measure
$\mu_{3}:=\nu_{2}=\frac{\sqrt{4-x^{2}}dx}{2\pi}.$
$(ii)$ Assume GST for $B(4)=E_{0}\times E_{1}\times E_{4}.$ Then the quantity $\delta_{p}(4)$ for $p$ running the set $\mathcal{P}_{1}:=\{p=4m+1\;\;is\;\ prime\}$ is equidistributed on [10,10][-10,10] with respect to the measure
$\mu_{1}:=\lambda_{cm}*\mu*\nu_{2},$
where $\mu*\lambda$ denotes the convolution of measures, and
$\lambda_{cm}(x)=\frac{1}{2}\nu_{1}(x/2)=\frac{\delta_{0}}{4}+\frac{dx}{2\pi\sqrt{16-x^{2}}},\mu(x)=\frac{1}{2}\nu_{2}(\frac{x}{2})=\frac{\sqrt{16-x^{2}}dx}{4\pi}.$
The proof follows directly from formulas (1),(8) and (6).
Therefore,
$\mu_{1}={h_{1}(x)dx},\;h_{1}=h_{11}+h_{12},\;\hbox{\rm for}$
$h_{11}:=\frac{\sqrt{16-s^{2}}*\sqrt{4-u^{2}}}{32\pi^{2}}\geq 0,\;h_{12}:=\frac{1}{16\pi^{3}\sqrt{16-v^{2}}}*\sqrt{16-s^{2}}*\sqrt{4-u^{2}}\geq 0.$
Remark. Note that the support of both $\lambda_{cm}$ and $\nu_{2}$ equals the whole interval [2,2][-2,2] thus
$\mathrm{supp}(\lambda_{cm})=\left[-4,4\right],\;\;\mathrm{supp}(\mu)=\left[-4,4\right],\;\;\mathrm{supp}(\nu_{2})=\left[-2,2\right],$
and, therefore,
$\mathrm{supp}(\mu_{3})=\left[-2,2\right],\quad\mathrm{supp}(\mu_{1})=\left[-10,10\right]\,.$
Indeed, if
$\mathrm{supp}(f)=[a_{1},b_{1}],\;\;\mathrm{supp}(g)=[a_{2},b_{2}],\;\;\mathrm{supp}(h)=[a_{3},b_{3}]$
then $\mathrm{supp}(f*g*h)=[a_{1}+a_{2}+a_{3},b_{1}+b_{2}+b_{3}].$
Remark. We have used that $E_{0},E_{1},E_{4}$ are not pairwise isogenous.
Clearly, $E_{0}$ has CM, and $E_{1},\bar{E}_{1},E_{4}$ have not, and thus $E_{0}$ is not isogenous (over $\bar{\mathbb{Q}}$) to $E_{1},\bar{E}_{1}$ or $E_{4};$ also, $E_{1},\bar{E}_{1}$ are not isogenous to $E_{4}$ , since $tr(E^{\prime}_{1}),tr(\bar{E}^{\prime}_{1})=\pm 2$ and $trE_{4}^{\prime}=-1$ for $E^{\prime}_{k}:=E_{k}\mod 5,$ and the traces are not equal up to sign. ∎
Note that GST($B(4)$) implies that the sequences $\{a_{0}(p)\},\{a_{1}(p)\}$ and $\{a_{4}(p)\}$ are statistically independent.
The condition in Theorem 4.1,(ii) is quite plausible, as mentioned in Section 3.
Looking at the support of $\mu_{1}$ and $\mu_{3}$, we obtain
Corollary 4.2.
Under the conditions of Theorem 4.1 for any $\varepsilon>0$ there exist $4$ primes $p_{1}=p_{1}(\varepsilon)\equiv 1\mod 4$, $p_{3}=p_{3}(\varepsilon)\equiv 3\mod 4,$ $p^{\prime}_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 4$, $p^{\prime}_{3}=p^{\prime}_{3}(\varepsilon)\equiv 3\mod 4,$ such that
$n_{p_{1}}(4)\geq\frac{p_{1}}{16}+\left(\frac{5}{8}-\varepsilon\right)\sqrt{p_{1}},\;\;n_{p^{\prime}_{1}}(4)\leq\frac{p^{\prime}_{1}}{16}-\left(\frac{5}{8}-\varepsilon\right)\sqrt{p^{\prime}_{1}},$
$n_{p_{3}}(4)\geq\frac{p_{3}}{16}+\left(\frac{1}{8}-\varepsilon\right)\sqrt{p_{3}},\;\;n_{p^{\prime}_{3}}(4)\geq\frac{p^{\prime}_{3}}{16}-\left(\frac{1}{8}-\varepsilon\right)\sqrt{p^{\prime}_{3}}.$
This corollary shows that the upper bounds for $n_{p}(4)-p/16$ are essentially tight. Note that the existence of $p_{3}$ and $p^{\prime}_{3}$ is unconditional.
4.3. Case $t=5$
For the next value $t=5$ one has $g(5)=17$ and $A(5)$ is (isogenous to) a product of 17 ECs; however, many of them are isomorphic or conjugate.
Indeed,
$A(5)=A(4)\times E_{5}\times E_{6}\times E_{7}\ldots\times E_{14}\times\mathrm{Jac}(C)$
for
$A(4)=E_{0}\times E_{1}\times E_{2}\times E_{3}\times E_{4}\equiv E_{0}^{2}\times E_{1}\times\bar{E}_{1}\times E_{4}$
with $E_{0},E_{1},E_{2},E_{3}$ and $E_{4}$ defined above;
$E_{5}:y^{2}=x(x+1)(x+4);\;E_{6}:y^{2}=x(x+2)(x+4);\;E_{7}:y^{2}=x(x+3)(x+4);$
$E_{8}:y^{2}=(x+1)(x+2)(x+4);\,E_{9}:y^{2}=(x+1)(x+3)(x+4);\,E_{10}:y^{2}=(x+2)(x+3)(x+4);\,$
$E_{11}:y^{2}=x(x+1)(x+2)(x+4);\,\;\;E_{12}:y^{2}=x(x+1)(x+3)(x+4);\;$
$E_{13}:y^{2}=x(x+2)(x+3)(x+4);\;\;E_{14}:y^{2}=(x+1)(x+2)(x+3)(x+4),\;$
$C:y^{2}=F(x)=x(x+1)(x+2)(x+3)(x+4),$
note that the curve $C$ admits a non-hyper-elliptic involution
$\sigma:(x,y)\mapsto\left(-\frac{2x+6}{x+2},\;\frac{2\sqrt{2}y}{(x+2)^{3}}\right)$
since $(x+2)^{6}F\big(\frac{2}{x+2}\big)=8F(x),$ and thus its trace can be calculated from the traces of a pair of elliptic curves, say, $E_{15}$ and $E_{16}$, where
$E_{15}:y^{2}=x^{3}+2\sqrt{2}x^{2}-9x-18\sqrt{2},$
$E_{16}:y^{2}=x^{3}+2x^{2}-9x/2-9,$
$E_{15},E_{16}$ being conjugate over $\mathbb{Q}(\sqrt{2}).$
One also notes numerous isomorphisms between those curves.
Namely,
$E_{0}$ is $\mathbb{Q}$-isomorphic to $E_{3},\;{{E_{6}}^{\prime}},\;E_{10};j(E_{0})=1728=2^{6}3^{3};$
$E_{4}$ is $\mathbb{Q}$-isomorphic to $E_{5},\;\bar{E}_{7},\;E_{14},\;j(E_{4})=35152/9=2^{4}3^{-2}13^{3};$
$E_{1}$ is $\mathbb{Q}$-isomorphic to $\bar{E_{2}},\;E_{8},\;\bar{E}_{9}\;E_{11},\;E_{13};\;j(E_{1})=21952/9=2^{6}3^{-2}7^{3};$
$J(E_{12})=1556068/81=2^{2}3^{-4}73^{3};$
$E_{15}\simeq_{\mathbb{Q}}\;{E_{16}^{\prime}},j(E_{15})=2744000/9=2^{6}3^{-2}5^{3}7^{3},$
where $\bar{E}$} is the $\mathbb{Q}(i)$-twist of $E,$ and ${E}^{\prime}$} is the $\mathbb{Q}(\sqrt{2})$-twist of $E.$
Therefore,
$A(5)=E_{0}^{3}\times E_{0}^{\prime}\times E_{1}^{4}\times\bar{E}_{1}^{2}\times E_{4}^{3}\times\bar{E}_{4}\times E_{12}\times E_{15}\times{E_{15}^{\prime}},$
where all elliptic curves in the product are pairwise non-isogenous.
Indeed, $E_{0},E_{1}$ and $E_{4}$ are pairly non-isogenous (over $\bar{\mathbb{Q}}$). For the pair $(E_{1},E_{12})$ it is sufficient to look at their reductions modulo 7 to see that $E^{\prime\prime}_{1}=E_{1}\mod 7$ is supersingular, $a_{1}(7)=0,$ whereas $E^{\prime\prime}_{12}=E_{12}\mod 7$ is not, $a_{12}(7)=-1.$ Thus $E_{1}^{\prime\prime}$ and $E^{\prime\prime}_{12}$ are not isogenous over $\bar{\mathbb{F}}_{7}$ and hence $E_{1},E_{12}$ are not isogenous over $\bar{\mathbb{Q}}$. The same for $(E_{1},E_{16})$ since $a_{16}(7)=5.$ For $(E_{4},E_{12}):a_{4}(11)=5,a_{12}(11)=-3;$ for $(E_{4},E_{16}):a_{4}(7)=1,a_{16}(7)=5;$ finally, $a_{12}(7)=-1,a_{16}(7)=5$ gives the same for $(E_{12},E_{16})$.
Admitting GST for $A(5),$ one gets a result parallel to Theorem 4.1 where the formulas are more lengthy and include more convolutions; in fact it is sufficient to admit GST for
$B(5):=E_{0}\times E_{1}\times E_{4}\times E_{12}\times E_{16},\;\dim B(5)=5.$
The results depend on $p$ modulo 8; for instance, in the simplest case $p=8m+7$ one has
$n_{p}(5)-\frac{p}{32}=\frac{a_{0}(p)}{16}+\frac{a_{1}(p)}{16}+\frac{a_{4}(p)}{16}+\frac{a_{12}(p)}{32},$
since $p$ is inert both in $\mathbb{Q}(\sqrt{-1})$ and $\mathbb{Q}(\sqrt{2})$ which gives the simplification of all other terms (as for $p=4k+3$ if $t=4$).
In general we get, using how $p$ splits in $\mathbb{Q}(\sqrt{-1})$ and $\mathbb{Q}(\sqrt{2})$:
Theorem 4.3.
$(i)$ Admit GST for
$B^{\prime}(5):=E_{0}\times E_{1}\times E_{4}\times E_{12},\;\dim B^{\prime}(5)=4.$
If $p=8m+7,$ then
$\big|\delta(p)\big|\leq\frac{7}{32}\;\;\;\hbox{ \rm{for}}\;\;\;\delta(p):=\frac{1}{2\sqrt{p}}\Big|n_{p}(5)-\frac{p}{32}\Big|.$
On the sequence $\mathcal{P}_{7}=\{p=8m+7\}$ the quantity $\delta_{7}(p)$ is equidistributed for the measure
$\mu_{7}(5)=\lambda_{cm}*\mu*\mu*\nu_{2}$
in the notation of Theorem 4.1 .
$(ii)$ Admit GST for $B^{\prime}(5).$
If $p=8m+5,$ then
$n_{p}(5)-\frac{p}{32}=\frac{a_{0}(p)}{16}+\frac{a_{1}(p)}{8}+\frac{3a_{4}(p)}{16}+\frac{a_{12}(p)}{32},$
and
$\big|\delta(p)\big|\leq\frac{13}{32}\;\;\;\hbox{ \rm{for}}\;\;\;\delta(p):=\frac{1}{2\sqrt{p}}\Big|n_{p}(5)-\frac{p}{32}\Big|.$
On the sequence $\mathcal{P}_{5}=\{p=8m+5\}$ the quantity $\delta(p)$ is equidistributed for the measure
$\mu_{5}(5)=\lambda_{cm}*\mu*\mu^{\prime}*\nu_{2}$
for $\mu^{\prime}:=\mu(x/3)/3.$
$(iii)$ Admit GST for $B(5).$
If $p=8m+3,$ then
$n_{p}(5)-\frac{p}{32}=\frac{a_{0}(p)}{8}+\frac{a_{1}(p)}{16}+\frac{a_{4}(p)}{16}+\frac{a_{12}(p)}{32}+\frac{a_{15}(p)}{16},$
and
$\big|\delta(p)\big|\leq\frac{11}{32}\;\;\;\hbox{ \rm{for}}\;\;\;\delta(p):=\frac{1}{2\sqrt{p}}\Big|n_{p}(5)-\frac{p}{32}\Big|.$
On the sequence $\mathcal{P^{\prime}}_{3}=\{p=8m+3\}$ the quantity $\delta(p)$ is equidistributed for the measure
$\mu_{3}(5)=\lambda^{\prime}_{cm}*\mu*\mu*\nu_{2}*\mu$
for $\lambda^{\prime}_{cm}:=\lambda_{cm}(x/2)/2.$
$(iv)$ Admit GST for $B(5).$
If $p=8m+1,$ then
$n_{p}(5)-\frac{p}{32}=\frac{a_{0}(p)}{8}+\frac{a_{1}(p)}{8}+\frac{3a_{4}(p)}{16}+\frac{a_{12}(p)}{32}+\frac{a_{15}(p)}{16},$
and
$\big|\delta(p)\big|\leq\frac{17}{32}\;\;\;\hbox{ \rm{for}}\;\;\;\delta(p):=\frac{1}{2\sqrt{p}}\Big|n_{p}(5)-\frac{p}{32}\Big|.$
On the sequence $\mathcal{P^{\prime}}_{1}=\{p=8m+1\}$ the quantity $\delta(p)$ is equidistributed for the measure
$\mu_{1}(5)=\lambda^{\prime}_{cm}*\mu*\mu^{\prime}*\nu_{2}*\mu.$
Corollary 4.4.
In the conditions of Theorem 4.3 .
$(i)$ One has the bound
$\big|\delta_{7}(p)\big|\leq\frac{17}{32}$
which is tight on $\mathcal{P^{\prime}}_{1}$.
$(ii)$ On $\mathcal{P^{\prime}}_{3}$ one has
$\big|\delta_{7}(p)\big|\leq\frac{13}{32},$
tight on $\mathcal{P^{\prime}}_{3}$.
$(iii)$ On $\mathcal{P}_{5}$ one has
$\big|\delta_{7}(p)\big|\leq\frac{11}{32},$
tight on $\mathcal{P}_{5}$.
$(iv)$ On $\mathcal{P}_{7}$ one has
$\big|\delta_{7}(p)\big|\leq\frac{7}{32},$
tight on $\mathcal{P}_{7}$.
5. Unconditional results
5.1. tt=4.
Note than that in [Jo] the GST(A) is proven, in particular, for any product $A=E\times F$ of two (non-isogenous) elliptic curves. Using this result, one can obtain an unconditional result similar to (but weaker than) Corollary 4.2. Indeed, one has
Lemma 5.1.
The sequences
$S_{0}=\{a_{0}(p)\},\;S_{1}=\{a_{1}(p)\},\;S_{4}=\{a_{4}(p)\}$
are pairwise statistically independent.
Indeed, the GST holds for $A_{01}=E_{0}\times E_{1},A_{04}=E_{0}\times E_{4},$ and $A_{14}=E_{1}\times E_{4}$ thanks to [Jo], which immediately implies that independence.
Proposition 5.2.
For any $\varepsilon>0$ there exist primes $p_{1},p_{1}^{\prime}\,$ with
$p_{1}=p_{1}(\varepsilon)\equiv 1\mod 4,\;p_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 4,$
such that
$n_{p_{1}}(4)\geq\frac{p_{1}}{16}+\left(\frac{3}{8}-\varepsilon\right)\sqrt{p_{1}},\;\;n_{p^{\prime}_{1}}(4)\leq\frac{p^{\prime}_{1}}{16}-\left(\frac{3}{8}-\varepsilon\right)\sqrt{p^{\prime}_{1}}.$
Proof. Fix $\varepsilon>0.$ Since $S_{0}$ and $S_{1}$ are independent, one can find $p_{1},p_{1}^{\prime}$ s.t.
$\frac{a_{0}(p_{1})}{\sqrt{p_{1}}}\geq 2-4\varepsilon,\;\frac{a_{1}(p_{1})}{\sqrt{p_{1}}}\geq 2-4\varepsilon,$
$\frac{a_{0}(p^{\prime}_{1})}{\sqrt{p_{1}}}\leq 4\varepsilon-2,\;\frac{a_{1}(p_{1})}{\sqrt{p^{\prime}_{1}}}\leq 4\varepsilon-2,$
and since $|{a_{4}(p_{1})}|\leq 2\sqrt{p_{1}}$, an immediate calculation implies the conclusion.∎
5.2. tt=5.
[…]
Proposition 5.3.
$(i)$ For any $\varepsilon>0$ there exist primes $p_{1},p_{1}^{\prime}\,$ with
$p_{1}=p_{1}(\varepsilon)\equiv 1\mod 8,\;p_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 8,$
such that
$n_{p_{1}}(5)\geq\frac{p_{1}}{32}+\left(\frac{1}{16}-\varepsilon\right)\sqrt{p_{1}},\;\;n_{p^{\prime}_{1}}(5)\leq\frac{p^{\prime}_{1}}{32}-\left(\frac{1}{16}-\varepsilon\right)\sqrt{p^{\prime}_{1}}.$
$(ii)$ For any $\varepsilon>0$ there exist primes $p_{3},p_{3}^{\prime}\,$ with
$p_{3}=p_{3}(\varepsilon)\equiv 3\mod 8,\;p^{\prime}_{3}=p_{3}^{\prime}(\varepsilon)\equiv 3\mod 8,$
such that
$n_{p_{3}}(5)\geq\frac{p_{3}}{32}+\left(\frac{3}{16}-\varepsilon\right)\sqrt{p_{3}},\;\;n_{p^{\prime}_{3}}(5)\leq\frac{p^{\prime}_{3}}{32}-\left(\frac{3}{16}-\varepsilon\right)\sqrt{p^{\prime}_{3}}.$
$(iii)$ For any $\varepsilon>0$ there exist primes $p_{5},p_{5}^{\prime}\,$ with
$p_{5}=p_{5}(\varepsilon)\equiv 5\mod 8,\;p^{\prime}_{5}=p_{5}^{\prime}(\varepsilon)\equiv 5\mod 8,$
such that
$n_{p_{5}}(5)\geq\frac{p_{5}}{32}+\left(\frac{1}{16}-\varepsilon\right)\sqrt{p_{5}},\;\;n_{p^{\prime}_{5}}(5)\leq\frac{p^{\prime}_{5}}{32}-\left(\frac{1}{16}-\varepsilon\right)\sqrt{p^{\prime}_{5}}.$
$(iv)$ For any $\varepsilon>0$ there exist primes $p_{7},p_{7}^{\prime}\,$ with
$p_{7}=p_{7}(\varepsilon)\equiv 7\mod 8,\;p^{\prime}_{7}=p_{7}^{\prime}(\varepsilon)\equiv 7\mod 8,$
such that
$n_{p_{7}}(5)\geq\frac{p_{7}}{32}+\left(\frac{1}{16}-\varepsilon\right)\sqrt{p_{7}},\;\;n_{p^{\prime}_{7}}(5)\leq\frac{p^{\prime}_{7}}{32}-\left(\frac{1}{16}-\varepsilon\right)\sqrt{p^{\prime}_{7}}.$
Indeed, e.g., in $(iv)$
$n_{p}(5)-\frac{p}{32}=\frac{a_{0}(p)}{16}+\frac{a_{1}(p)}{16}+\frac{a_{4}(p)}{16}+\frac{a_{12}(p)}{32},$
and one can choose first two terms very close to $\sqrt{p}/8$ or to $\sqrt{p}/8,$ by independence of $a_{0}(p)$} and $a_{1}(p)$}. Similar arguments work for all other cases.